What is the prime factorization of 100
Answer by Edwin McCravy(20086)

Make a factor tree:

        100 <- The smallest prime that goes into 100 is 2, and 100 = 2×50
         /\
        2  50 <- The smallest prime that goes into 50 is 2, and 50 = 2×25
           /\
          2  25 <- The smallest prime that goes into 25 is 5, and 25 = 5×5
             /\
            5  5 <- 5 is prime so you're done.

100 = 2×2×5×5 = 2²×5²

Edwin